The table below shows the probability distribution of the random variable X.
a. Find the mean of the random variable.
b. Obtain the standard deviation sigma of the random variable.
x P(X=x)
1 0.1
2 0.4
3 0.5

Answer by Boreal(15235)

mean or expected value is sum of x*p(x)
that is 0.1+0.8+1.5 or 2.4
sd
first, [x-(E(X))^2 * p(x)]
-1.4^2*0.1=.196
-0.4^2*0.4=0.064
0.6^2*0.5=0.18
sum is 0.44
Take the square root = 0.66 is sd